OpenStudy (anonymous):

find an equation for the line through (-4,-8) horizantal

OpenStudy (yrelhan4):

horizontal means slope is 0.. basic equation os a line.. y-y1/x-x1=m.. just substitute the values of x1,y1 that you are given and m as i told you.. and you have your answer..

OpenStudy (anonymous):

ok i got -4 is that correct?

OpenStudy (yrelhan4):

-4? you need to find the equation of the line.. [y-(-8)]/[x-(-4)]=0 ==> y+8=0 ==> y=-8.. thats your answer

OpenStudy (anonymous):

ohhh ok thank you i appriciate it:)

OpenStudy (yrelhan4):

You're welcome.